• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Calarco

    Log entry time 16:03:10 on November 4,2000

    Entry number 51470

    keyword=Shift summary

    0810: Gary Dezern here to work on right HRS collimator. MCC guy says that
    we need "official" Radcon permission to work in that area .... waiting for that.
    0840: Problem solved. There is a Kepco supply underneath the spectrometer that
    supplies power to the collimator motor. The supply had tripped. We reset it and
    moved the 6msr collimator into place from the medm window on the hall PC.
    The plan is to check the coincidence rate with the 6msr collimator on the electron
    arm (HRS-right) and then give the beam over to the Moeller measurement to be
    0910: Machine problems since we came out of the hall .... waiting for beam.
    1150: First sniff of CW beam, I=2.5 microA, checked that raster is on and looks OK.
    1230: Have established that coincidence rate is as expected, 90 Hz with open
    collimator left, 6msr right (see run 1059) and have now installed sieve collimator
    1240: started run 1060: LH2 open left, sieve right
    1410: end of run 1060 failed, reset CODA; changed collimator to open left, open right
    1420: start of run 1061 failed, so we killed CODA and restarted ... new run is 1062
    1430: for past runs since 1200 we note that DAQ deadtime is 0% with rates as high
    as 450 Hz
    1516: started straw efficiency run 1062
    1528: run 1063 PS 1/10/1/20/1 to check DT: event rate now 1600, still no deadtime.
    1532: set PS to 1/20/1/2000/1 and started run 1064; deadtime still reads 0; Bob
    Michaels working on it.